# Dynamics due to competitive flip cycles in active Potts models

### v1

@@ Abstract
Nonequilibrium spatiotemporal patterns have been extensively studied. However, a single oscillator or cyclic loop of states is typically employed at each site in theories and simulations. Here, we investigated how competition among multiple identical cyclic loops at each site alters patterns. We simulated active Potts models with standard Potts interactions between neighboring sites in two-dimensional square lattices. When multiple three-state cycles exist in state flips, such as in octahedral and square-antiprism networks, all types of spiral waves comprising the three states are formed simultaneously at high flip energies. However, at lower energies, only one or a few types emerge and switch stochastically into different types. At even lower energies, cyclic changes in single-state dominant homogeneous phases emerge [homogeneous cycling (HC) mode]. At intermediate flip energies, the spiral wave and HC modes temporally coexist in small systems but do not switch between each other in large systems. Conversely, when multiple four-state cycles exist in six-state and cubic networks, one state remains dominant for the entire range of flip energies, whereas the other states occasionally form domains at intermediate flip energies. Therefore, the number of spatially coexisting states can be controlled using flip networks and energies.
@@ I Introduction
The models and methods are described in Sec. II. Simulation results for the three- and four-state-cycle networks are presented and discussed in Secs. III and IV, respectively. The dynamics for the single three-state and four-state cycles are briefly reviewed in Secs. III.1 and IV.1, respectively. The results of the four-state models with two three-state cycles are described in Sec. III.2. A theoretical analysis at the absence of contact interactions is presented in Sec. III.2.1. The results of the six- and eight-state models with the octahedral and square-antiprism networks are described in Secs. III.3, and III.4, respectively. The results of six- and eight-state models with two four-state cycles and a cubic network are described in Secs. IV.2, and IV.3, respectively. Finally, a summary is presented in Sec. V.
@@ III.2.1 Homogeneous Mixed State in the Absence of Interactions
Figure 3: Active Potts model with two three-state cycles [Fig. 1(a)] at $L=256$. (a) Snapshots at $h=0.7$, $1.5$, and 22 (from left to right). The light yellow, green, blue, and red sites (light to dark in grayscale) represent $s=0$, 11, 22, and $3$, respectively. (b)–(e) Time development of the number of each state at (b) $h=2$, (c) $h=1.5$, and (d),(e) $h=0.7$. Two three-state spiral waves ($s=0\to 1\to 2\to 0$, and $s=3$ instead of $s=1$) spatially coexist (W4) at $h\gtrsim 0.9$, and the $s=0$ state becomes dominant at high $h$. The three-state spiral waves (W3) and homogeneous cycling (HC) are obtained at $h=0.7$, depending on the initial state. In W3, the two types of waves temporally coexist.
Figure 4: Dependence on the flip energy $h$ in the active Potts model with two three-state cycles [Fig. 1(a)]. (a)–(b) Mean number density $\langle N_{s}\rangle/N$ at (a) $L=256$ and (b) $L=128$. The blue down-pointing triangles in (a) represent $s=2$ in the HC mode. (c) Time fractions $p_{\mathrm{phase}}$ of phases at $L=256$ (solid lines) and $L=128$ (dashed lines). The red down-pointing triangles represent the ratio of the single-state dominant phase in the HC mode at $L=256$, whereas the ratios of multiple-state coexisting phases in the HC mode are omitted for clarity. The bidirectional arrows at the top of (a) represent the ranges of three modes at $L=256$.
@@ III.2.2 Dynamics in the Absence of Diagonal Flips
We consider herein the dynamics under contact interactions ($J=2$) for two three-state cycles [Fig. 1(a)]. At low $h$, cyclic changes in homogeneous phases occur via two routes ($s=0\to 1\to 2\to 0$ and $s=0\to 3\to 2\to 0$) i.e., the HC mode emerges [the numbers of states $N_{s}/N\approx 1$ are switched in Fig. 3(e)]. At high $h$, the domains of all four states spatially coexist and two types of spiral waves ($s=0\to 1\to 2\to 0$ and $s=0\to 3\to 2\to 0$) are steadily formed (W4) [see Fig. 3(a)–(c)]. At intermediate $h$, one type of spiral wave is dominantly formed (W3), and the switch to the other type occurs stochastically [the positions of the red diamonds and yellow triangles are switched several times in Fig. 3(d)]. In the spiral waves shown in the left snapshot of Fig. 3(a) and Movie S1, small domains ($s=1$) often appear due to nucleation, but most of them eventually disappear owing to fluctuations. With increasing $h$, the HC and W3 modes begin to coexist temporally, and the ratio of the W3 mode gradually increases at $L=128$. Conversely, the switching between two modes does not occur in available simulation periods ($\sim 10^{8}$) at $L=256$; the HC and W3 modes are obtained when the uniform and random distributions of states are respectively used as initial states for $0.3\lesssim h\lesssim 0.7$.
@@
Figure 6: Active Potts model with the octahedral network [Fig. 1(c)] at $L=512$. (a) Snapshots at $h=0.6$ (left) and $h=1$ (right). The light yellow, light green, cyan, blue, magenta, and red sites (light to dark in grayscale) represent $s=0$, 11, 22, 33, 44, and $5$, respectively. (b)–(d) Time development of the number of each state at (b) $h=1$ and (c),(d) $h=0.6$. Eight three-state spiral waves spatially coexist (W8) at $h=1$. The three-state spiral waves (W3) and homogeneous cycling (HC) are obtained at $h=0.6$, depending on the initial state. In W3, the eight types of waves temporally coexist.
@@ III.2.3 Dynamics in the Presence of Diagonal Flips
When diagonal flips between the $s=1$ and $3$ states are added [Fig. 1(b)], the number density of the $s=3$ state decreases with increasing $h_{\mathrm{a}}$, as in the case in which there are no contact interactions [compare Figs. 5(b) and 2]. However, the densities of the $s=2$ and $s=0$ states also change in the presence of diagonal flips, unlike its values in their absence. For $h=2$, the $s=2$ density monotonically increases with increasing $h_{\mathrm{a}}$, whereas a minimum appears at $h_{\mathrm{a}}/h\simeq 0.4$ for $h=1$ owing to a decrease in the time fraction of the HC mode. The flips from the $s=3$ to $s=1$ states yield small clusters of $s=1$ and $3$ on the domain boundary of $s=0$, so that the flips from the $s=0$ to $s=1$ or $3$ are enhanced, resulting in a decrease in $\langle N_{0}\rangle/N$ (see Fig. 5).
@@ III.4 Eight-state Potts Model with Square-Antiprism Network
In all three networks, spiral waves of a single three-state cycle or multiple cycles emerge. Therefore, the spiral wave is a robust mode for the three-state cycles. The number of cycles and states can vary according to the network and flip energy choices.
@@ IV.1 Single Four-state Cycle
Before discussing the competition among multiple four-state cycles, we briefly describe the dynamics induced by the single four-state cycle in the four-state active standard Potts model under cyclic symmetry [31]. At low $h$ ($h>0$), the four homogeneous phases change cyclically as $s=0\to 1\to 2\to 3\to 0$ through nucleation and growth (HC mode), as in the three-state model. At high $h$, the four states spatially coexist, but the waves do not have spiral centers (W4 mode) under the interaction of the standard Potts model [33]. The domain boundaries between neighboring states propagate ballistically, but the boundaries between non-neighboring states move slowly (driven by diffusion). Hence, the existence of non-flip diagonal states alters the dynamics. The transition between the HC and W4 modes occurs via temporal coexistence, even at $L=512$. At intermediate flip energies, the domains of the diagonal states often remain for long periods, slowly shrinking to reduce the interfacial energy.
@@ IV.3 Eight-state Potts Model with cubic Network
The even-numbered states dominantly covered the lattice, as the $s=3$ state in the case of the six-state model described in Sec. IV.2 (see Figs. 13 and 14). The HC mode appears at low $h$ [see Fig. 13(c)], and diagonal domains are formed at intermediate and high $h$ [see Fig. 13(a)–(c) and Movie S6]. In contrast to the six-state model, these domains are larger and have longer lifetimes. This is due to the existence of more diagonal states (e.g., $s=2$, 44, 66, and $7$ for $s=0$), and three of them exhibit the same stability ($s=2$, 44, and $6$ for $s=0$). This domain formation also occurs in the HC mode [see Fig. 13(d)].
In both networks, the competition among the four-state cycles suppresses the traveling waves. The single-state dominant phases are stabilized, and the domains of the diagonal states often coexist.
@@ V Summary
Regarding the competition among the four-state cycles, two networks are examined: a six-state network of two cycles and an eight-state cubic network. Single-state dominant phases occupy the systems most of the time for all simulation conditions, so that the steady traveling waves are suppressed by competition. At intermediate $h$, diagonal states often form domains but subsequently shrink slowly. This is due to the absence of direct flips between the diagonal states. In this study, we employed the contact interactions for the standard Potts model. Spiral waves and other forms of dynamics are likely to emerge when different contact interactions are used in active Potts models of four or more states.

### v2

@@ Abstract
Nonequilibrium spatiotemporal patterns have been extensively studied. However, a single oscillator or cyclic loop of states is typically employed at each site in theories and simulations. Here, we investigate how competition among multiple identical cyclic loops at each site alters patterns. We simulate active Potts models with standard Potts interactions between neighboring sites in two-dimensional square lattices. When multiple three-state cycles exist in state flips, such as in octahedral and square-antiprism networks, all types of spiral waves comprising the three states are formed simultaneously at high flip energies. However, at lower energies, only one or a few types emerge and switch stochastically into different types. At even lower energies, cyclic changes in single-state dominant homogeneous phases emerge [homogeneous cycling (HC) mode]. At intermediate flip energies, the spiral wave and HC modes temporally coexist in small systems but do not switch between each other in large systems. Conversely, when multiple four-state cycles exist in six-state and cubic networks, one state remains dominant for the entire range of flip energies, whereas the other states occasionally form domains at intermediate flip energies. Therefore, the number of spatially coexisting states can be controlled using flip networks and energies.
@@ I Introduction
The models and methods are described in Sec. II. Simulation results for the three- and four-state-cycle networks are presented and discussed in Secs. III and IV, respectively. The dynamics for the single three-state and four-state cycles are briefly reviewed in Secs. III.1 and IV.1, respectively. The results of the four-state models with two three-state cycles are described in Sec. III.2. A theoretical analysis at the absence of contact interactions is presented in Sec. III.2.1. The results of the six- and eight-state models with the octahedral and square-antiprism networks are described in Secs. III.3 and III.4, respectively. The results of six- and eight-state models with two four-state cycles and a cubic network are described in Secs. IV.2 and IV.3, respectively. Finally, a summary is presented in Sec. V.
@@ III.2.1 Homogeneous Mixed State in the Absence of Interactions
Figure 3: Active Potts model with two three-state cycles without diagonal flips [Fig. 1(a)] at $L=256$. (a) Snapshots at $h=0.7$, $1.5$, and 22 (from left to right). The light yellow, green, blue, and red sites (light to dark in grayscale) represent $s=0$, 11, 22, and $3$, respectively. (b)–(e) Time development of the number of each state at (b) $h=2$, (c) $h=1.5$, and (d),(e) $h=0.7$. Two three-state spiral waves ($s=0\to 1\to 2\to 0$, and $s=3$ instead of $s=1$) spatially coexist (W4) at $h\gtrsim 0.9$, and the $s=0$ state becomes dominant at high $h$. The three-state spiral waves (W3) and homogeneous cycling (HC) are obtained at $h=0.7$, depending on the initial state. In W3, the two types of waves temporally coexist as shown in (d).
Figure 4: Dependence on the flip energy $h$ in the active Potts model with two three-state cycles without diagonal flips [Fig. 1(a)]. (a)–(b) Mean number density $\langle N_{s}\rangle/N$ at (a) $L=256$ and (b) $L=128$. The blue down-pointing triangles in (a) represent $s=2$ in the HC mode. (c) Time fractions $p_{\mathrm{phase}}$ of phases at $L=256$ (solid lines) and $L=128$ (dashed lines). The red triangles represent the ratio of the single-state dominant phase in the HC mode at $L=256$, whereas the ratios of multiple-state coexisting phases in the HC mode are omitted for clarity. The bidirectional arrows at the top of (a) represent the ranges of three modes at $L=256$.
@@ III.2.2 Dynamics in the Absence of Diagonal Flips
We consider herein the dynamics under contact interactions ($J=2$) for two three-state cycles [Fig. 1(a)]. At low $h$, cyclic changes in homogeneous phases occur via two routes ($s=0\to 1\to 2\to 0$ and $s=0\to 3\to 2\to 0$) i.e., the HC mode emerges [the numbers of states $N_{s}/N\approx 1$ are switched in Fig. 3(e)]. At high $h$, the domains of all four states spatially coexist and two types of spiral waves ($s=0\to 1\to 2\to 0$ and $s=0\to 3\to 2\to 0$) are steadily formed (W4) [see Fig. 3(b),(c) and the right two snapshots in Fig. 3(a)]. At intermediate $h$, one type of spiral wave is dominantly formed (W3), and the switch to the other type occurs stochastically [the positions of the red diamonds and green triangles are switched several times in Fig. 3(d)]. In the spiral waves shown in the left snapshot of Fig. 3(a) and Movie S1, small domains ($s=1$) often appear due to nucleation, but most of them eventually disappear owing to fluctuations. With increasing $h$, the HC and W3 modes begin to coexist temporally, and the ratio of the W3 mode gradually increases at $L=128$. Conversely, the switching between two modes does not occur in available simulation periods ($\sim 10^{8}$) at $L=256$; the HC and W3 modes are obtained when the uniform and random distributions of states are respectively used as initial states for $0.3\lesssim h\lesssim 0.7$.
@@
Figure 6: Active Potts model with the octahedral network [Fig. 1(c)] at $L=512$. (a) Snapshots at $h=0.6$ (left) and $h=1$ (right). The light yellow, light green, cyan, blue, magenta, and red sites (light to dark in grayscale) represent $s=0$, 11, 22, 33, 44, and $5$, respectively. (b)–(d) Time development of the number of each state at (b) $h=1$ and (c),(d) $h=0.6$. All types of three-state spiral waves spatially coexist (W6) at $h=1$. The three-state spiral waves (W3) and homogeneous cycling (HC) are obtained at $h=0.6$, depending on the initial state, [(c) and (d), respectively]. In W3, the eight types of waves temporally coexist.
@@ III.2.3 Dynamics in the Presence of Diagonal Flips
When diagonal flips between the $s=1$ and $3$ states are added [Fig. 1(b)], the number density of the $s=3$ state decreases with increasing $h_{\mathrm{a}}$, as in the case in which there are no contact interactions [compare Figs. 5(b) and 2]. However, the densities of the $s=2$ and $s=0$ states also change in the presence of diagonal flips, unlike the values in their absence. For $h=2$, the $s=2$ density monotonically increases with increasing $h_{\mathrm{a}}$, whereas a minimum appears at $h_{\mathrm{a}}/h\simeq 0.4$ for $h=1$ owing to a decrease in the time fraction of the HC mode. The flips from the $s=3$ to $s=1$ states yield small clusters of $s=1$ and $3$ on the domain boundary of $s=0$, so that the flips from the $s=0$ to $s=1$ or $3$ are enhanced, resulting in a decrease in $\langle N_{0}\rangle/N$ (see Fig. 5).
@@ III.4 Eight-state Potts Model with Square-Antiprism Network
In all three networks, spiral waves of a single three-state cycle or multiple cycles emerge. Therefore, the spiral wave is a robust mode for the three-state cycles. The number of cycles and states can vary according to the network and flip energy choices. The competition among the cycles reduce the number of spatially coexisting states at intermediate flip energies.
@@ IV.1 Single Four-state Cycle
Before discussing the competition among multiple four-state cycles, we briefly describe the dynamics induced by the single four-state cycle in the four-state active standard Potts model under cyclic symmetry [31]. At low $h$ ($h>0$), the four homogeneous phases change cyclically as $s=0\to 1\to 2\to 3\to 0$ through nucleation and growth (HC mode), as in the three-state model. At high $h$, the four states spatially coexist, but the waves do not have spiral centers (W4 mode) under the interaction of the standard Potts model [31, 33]. The domain boundaries between neighboring states propagate ballistically, but the boundaries between non-neighboring states move slowly (driven by diffusion). Hence, the existence of non-flip diagonal states alters the dynamics. The transition between the HC and W4 modes occurs via temporal coexistence, even at $L=512$. At intermediate flip energies, the domains of the diagonal states often remain for long periods, slowly shrinking to reduce the interfacial energy.
@@ IV.3 Eight-state Potts Model with Cubic Network
The even-numbered states dominantly covered the lattice, as the $s=3$ state in the case of the six-state model described in Sec. IV.2 (see Figs. 13 and 14). The HC mode appears at low $h$ [see Fig. 13(d)], and diagonal domains are formed at intermediate and high $h$ [see Fig. 13(a)–(c) and Movie S6]. In contrast to the six-state model, these domains are larger and have longer lifetimes. This is due to the existence of more diagonal states (e.g., $s=2$, 44, 66, and $7$ for $s=0$), and three of them exhibit the same stability ($s=2$, 44, and $6$ for $s=0$). This domain formation also occurs in the HC mode [see Fig. 13(d)].
In both networks, the competition among the four-state cycles suppresses the traveling waves. The single-state dominant phases are stabilized, and the domains of the nonflip (diagonal) states often coexist. Therefore, the existence of nonflip states in the flip cycles alters the dynamics.
@@ V Summary
Regarding the competition among the four-state cycles, two networks were examined: a six-state network of two cycles and an eight-state cubic network. Single-state dominant phases occupy the systems most of the time for all simulation conditions, so that the steady traveling waves are suppressed by competition. At intermediate $h$, diagonal states often form domains but subsequently shrink slowly. This is due to the absence of direct flips between the diagonal states in the four-state cycles.
Let us discuss how the flip-cycle competition changes spatiotemporal patterns. In the competitive three-state cycles, the single-cycle spiral waves of three states are obtained at small networks (the four- and six-state networks) with intermediate flip energies. Emerged waves suppress the growth of neighboring states so that spiral waves comprising other states do not spatially coexist. However, cycles separated in the flip network can coexist at large networks, and waves of different cycles can simultaneously appear. The number of cycles in waves increases in a larger network. Conversely, the competition among multiple four-state cycles suppresses all waves, leading to single-state dominant phases. Therefore, the competition can be used to reduce the wave types or erase the waves using three- or four-state cycles, respectively.
In this study, we employed the contact interactions for the standard Potts model. For a single cycle of four or more states, waves do not exhibit spiral shapes in the standard Potts model. However, spiral waves are formed when the contacts of nonflip states are repulsive, and factorized-symmetry modes also emerge for cyclic loops of factorizable numbers of states (e.g., $6=3\times 2$[33]. Therefore, the competition among spiral waves and other forms of dynamics can be explored when different contact interactions are used in active Potts models with multiple cycles of four or more states.